Janet has been able to divide her own age into her grandmother’s age evenly for the past six years. How old are they today?
Answer by VFBundy(438)

6 and 66

66/6 = 11
65/5 = 13
64/4 = 16
63/3 = 21
62/2 = 31
61/1 = 61